Two short dipoles each of the dipole moment p are placed at the origin. The dipole moment of one dipole is along the x-axis while that of the other is along the y-axis. The electric field at a point (a,0) is given by

Grade:12

1 Answers

Samyak Jain
5 years ago
Electric field at a point on the axis distant ‘a’ from a short dipole of dipole moment p is 2Kp/a3
(here bold letter denotes vector and K is electrostatic constant)
i.e. here, electric field at point (a,0) due to short dipole along x-axis is 2Kp/a3 i.
Electric field at an equatorial point (point on perpendicular bisector of dipole) distant ‘a’ from a short dipole of dipole moment p is – Kp/a3
(here too bold letter denotes vector)
i.e. here, magnitude of electric field at point (a,0) due to short dipole along y-axis is – Kp/a3 j.
\therefore Resultant electric field at point (a.0) is 2Kp/a3 i – Kp/a3 j.
Its magnitude is |2Kp/a3 i – Kp/a3 j| = \sqrt{(2Kp/a^3)^2 + (-Kp/a^3)^2}  =  (Kp/a3)\sqrt{2^2 + 1}
                                                    = \sqrt{5}Kp/a3.